OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4>

OpenStudy (slaaibak):

pretty sure the formula is cos^-1(u'v/(||U||*||v||))

OpenStudy (anonymous):

can you show me how to plug in the numbers @slaaibak

OpenStudy (slaaibak):

u'v = <6, -1>'<7,-4> = 6 * 7 + (-4)*(-1) = 42 + 4 = 46 ||v|| = sqrt(v'v) use same method as above for v'v